OpenStudy (anonymous):

quotient of x over 6y and (6xy)^2 over 5xy^3

OpenStudy (anonymous):

so is it \[\frac{ x }{ 6y }+ \frac{ (6xy)^{2} }{ 5xy ^{3} }\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok give me one sec. :)

OpenStudy (anonymous):

would it be \[\frac{ 36x ^{2} y}{ 11y ^{3} }\]

OpenStudy (anonymous):

i do homework online so i put your answer and its saying its wrong :d

OpenStudy (anonymous):

oh i'm sorry.... would you know why it would be wrong? :( (sorry)

OpenStudy (anonymous):

no, i dont know ive been stock in this problem for almost 20min :/

OpenStudy (anonymous):

ohhh...well i mean i don't know what would make it wrong because when you work out the problem you would first solve the exponent... then add like bases... though you can't add x+ x^2 because they are different... same thing with the denominator then you would simplify.... i'm sorry i couldn't help you... :(

OpenStudy (anonymous):

no problem, thank you very much tho :3

OpenStudy (anonymous):

no problem i'm just upset i couldn't help you... good luck with the problem though! :/